The hypothenuse of a right triangle measures 25 cms. A leg of the triangle measures 15 cms. What is the area of the triangle? please show work/explaition thank you so much :D
--sandy

Answer by Osran_Shri(18)

Data Given:
Right Angle Traingle
Hypotenuse = 25 cm
One leg = 15 cm
Let C=Hypotenuse = 25 cm
One leg = A = 15 cm
Other leg = B ( to be claculated before calculating area)
In a right angle traingle , according to Pythagorean theorem,
C^2 = A^2 + B^2
25^2 = 15^2 + B^2
625 = 225 + B^2
625 - 225 = B^2
400 = B^2
Therefore B = Sqrt of 400 = 20
Area of a Right angle triangle = (One leg * Other leg) / 2
That is Area = (A * B) / 2 = (15 * 20)/2 = 300/2 = 150 Square cm
Please contact me if you require any more queries / problems
osran_con@yahoo.co.in